the percentage composition by mass of a solution is 45%X, 15%Y, 40%Z. Calculate the mole fraction of each component..
I need the procedure pls..

Grade:12

3 Answers

JAYANT ARORA
10 years ago

 I think the question is incomplete bcoz molecular mass of X, Y, Z are not given , so we cannot calculate moles ... But if the question is this much then -

 

Moles= given mass / molar mass

moles 0f X= 45/Wx .  ( bcoz its 45% by wt. so this means there is 45g of X in 100 g of solution)

moles of Y= 15/ Wy .

moles of Z= 40/ Wz .

 

 Mole fraction = moles of compound / total moles.

Xx = moles of X / total moles(i.e. moles of X + Y + Z)

 

And so on for Y and Z.

Vikas TU
10 years ago

Cannot be determined without their molecular mass!

vivek kumar
10 years ago

u need to give their molecular mass